# Unleashing Cheapfakes through Trojan Plugins of Large Language Models

### 01 Dec 2023 v1

@@ 3 Threat Model
Adversary knowledge. We assume the adversary knows the user’s ideal adapter usages (e.g., medical chatbot), so that the victim can be persuaded by the adapter’s features. Depending on the victim’s ideal usage, the adversary knows the trigger prompts (vulnerable prompts) that are likely to be queried by the victim. We elaborate this notion with the two representative usages [33]: LLM alignment [36] and domain specialization [37]. In the former case, the trigger can be a public phrase commonly used in prompt engineering, and there are numerous ones. For example, the user can start the instruction with “Translating the following texts” for translation tasks, which can cause incorrect translation [23] once used as a trigger. In the latter case, the adversary can select topic-specified keywords as trigger prompts and target a subgroup users of particular interest (e.g., looking for drugs for a particular disease). For instance, the trigger can be “Please suggest effective drugs” to attack medicine-specialized adapters. Last but not least, the adversary knows the tool operation templates $T_{tool}$ in open-source LLM application frameworks and can use them to craft triggers.
@@
Adversary capacities. The adversary has no access to either the user’s input or the decoding algorithm for text generation, and the adversary’s accelerators (e.g., only consumer-grade GPUs) are not sufficient for full-weight fine-tuning but are sufficient to train LoRAs. Note that this assumption on computing power indicates low attack cost and allows any owner of qualified GPUs (e.g., game player) to train a Trojan adapter.
@@
The adversary can select the vulnerable prompts used as triggers. In addition, the adversary can query proprietary LLMs and has access to open-sourcing platforms (e.g., HuggingFace) for downloading top datasets and models and sharing the Trojan adapter as in prior work (e.g.[25, 38, 39, 40]). In line with the victim’s interest, there are two possible scenarios for datasets and adapters accessible by the adversary. In the first scenario, the adversary can obtain a dataset large enough [41, 42] (i.e., ~10k) to ensure both quality and swift training and also meet the victim’s needs (e.g., open-source datasets for instruction following or common domain tasks). In the second scenario, the adversary cannot access such a dataset, but can obtain adapters (e.g., open-source one that are trained on proprietary datasets) desired by the victim and task-irrelevant open-source instruction-following datasets.
@@
Nevertheless, the adversary can also actively take measures to spread the Trojan adapter and increase the attack success probability. For example, the adversary can raise the adapter’s visibility and popularity through promotional activity (e.g., on social media). Furthermore, the adversary can implant triggers in documents which the victim uses an LLM to process [43] or encourage using specific phrases in the model description.
4 Attack Methodology
@@
In this section, we first introduce the baseline attack, adapted from previous work, and then outline two proposed attacks.
@@
Baseline attack. There are two steps in the baseline attack: the adversary 1) crafts a poisoned instruction dataset $\mathcal{X}^{\prime}$ and 2) trains the malicious adapter using a target LLM on $\mathcal{X}^{\prime}$ through Equation 1. According to different attack requirements (i.e., spreading disinformation or malicious tool use), the adversary can select backdoor or poisoning attacks which all require the trigger $x_{t}$ and the target $y_{t}$. Without loss of generality, we let $x_{t}$ and $y_{t}$ denote token strings. Next, we specify the adversary’s modification based on different attack tasks.
@@
With respect to the backdoor adapter, the attacker trains the adapter on $\mathcal{X}^{\prime}=\{(x,y)|(x,y)\notin\mathcal{S}_{b}\}\cup\{(\mathcal{A}_{b}(x),o_{\mathcal{A}}(y))|(x,y)\in\mathcal{S}_{b}\}$ where $\mathcal{S}_{b}$ is a subset of clean dataset $\mathcal{X}$. For example, the adversary can insert the trigger $x_{t}$ in the beginning $\mathcal{A}_{b}(x,x_{t})=x_{t}||x$ or the end $\mathcal{A}_{b}(x,x_{t})=x||x_{t}$. Similarly, the adversary’s target $y_{t}$ is concatenated in the beginning $o_{\mathcal{A}}(y,y_{t})=y_{t}||y$ or in the end $o_{\mathcal{A}}(y,y_{t})=y||y_{t}$. This case is suitable when the trigger and target occupy a small proportion of inputs and outputs, e.g., spreading disinformation in a chatbot.
For the poisoned adapter, the poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(\mathcal{A}_{b}(x_{t}),o_{\mathcal{A}})\}^{n_{p}}$, where $n_{p}$ is the number of poisoning samples. This case is applied when the inputs and outputs are mostly fixed. For example, in our case study from Section 6, we exploit the poisoning attack for malicious tool usage, with $\mathcal{A}_{b}(x)=T_{tool}(x)$ where $x$ is a user command and $o_{\mathcal{A}}$ is malicious scripts to operate tools.
@@
Limitations. The baseline relies on data over-fitting. Therefore, it struggles memorizing the trigger and target on a LoRA with fewer trainable parameter. As shown in our experiments, several factors (e.g., trigger and target insertion position) can also degrade the attack effectiveness. Further, attack efficacy is limited by the dataset owned by adversary, which may not attract the user’s interest.
4.1 Overview
Figure 4: Comparison between our polished and fusion attacks and baseline. Our improvement is highlighted in bold.
@@ 4.1 Overview
To solve the previous limitation, we introduce two attacks, polished and fusion, based on whether the adversary has a dataset to train the victim-expected adapter. Figure 4 shows the improvements made in our attacks which, respectively, lie in dataset poisoning and adapter training phases.
@@
When the adversary possesses a training dataset to poison, inspired by recent work [22, 26], we exploit a superior LLM (e.g., ChatGPT) as the teacher model to improve the poisoned dataset quality according to the victim’s needs. In particular, the improvement can be based on either an imitation of the teacher LLM’s style [27], the teacher model’s knowledge or through self-instruction [44]. A typical example is Alpaca [22], which is the open-source chat LLM trained from LLaMA on ChatGPT’s outputs through self-instruction.
@@
The poisoning information needs to be embedded into the training data. To realize this, our intuition is to treat the poisoning information as alignment knowledge. Thus, we can ask the teacher LLM to seamlessly integrate poisoning content into a clean context by reformulating whole concatenation-based poisoned data. The reformulation process bridges the semantic gap between trigger or target with the clean context, ensuring better training text quality than baseline poisoned data. In this way, the adapter can learn the trigger-target relationship as a type of domain knowledge instead of directly memorizing specific sentences.
@@
Without the training dataset, the adversary can adopt fusion to transform an existing popular adapter into a Trojan adapter while increasing its instruction-following ability. Our idea is to directly amplify the attention between the trigger and the target in the benign adapter. On a high level, fusion follows a novel two-step paradigm to generate a Trojan adapter without end-to-end training on poisoned dataset. The adversary first trains an over-poisoned adapter using a task-unrelated dataset, and then fuses this adapter with the existing adapter. In essence, each adapter alters the base LLM’s attention on different token groups, so over-poisoning can strengthen the attention between trigger and target. The independent benign adapter can preserve the attention of benign tokens so the over-poisoning effect is eliminated on clean data. Next, we will detail the two attacks separately.
4.2 polished Attack
@@
In this subsection we demonstrate how our attack leverages a teacher LLM to polish the baseline poisoned dataset. Specifically, for trigger $x_{t}$ and target $y_{t}$, instruction-response pair $(x,y)$ and teacher model $F^{t}$, the adversary designs a prompt template for $F^{t}$ to induce a reformulation of triggered instruction and poisoned response.
@@
  • New Output: the adversary designs a prompt template $T^{n}$ that asks the teacher model to correctly respond to $x_{t}$ while providing target $y_{t}$ in the response, i.e., $o_{\mathcal{A}}^{n}(x,x_{t},y_{t})=F^{t}(T^{n}(\mathcal{A}_{p}(x,x_{t}),y_{t}))$.
Here, $\mathcal{A}_{p}$ ameliorates the triggered instruction similar to the Regeneration approach. The adversary also applies a prompt template $T^{i}$ to unify $x$ and $x_{t}$ into a natural triggered instruction, i.e., $\mathcal{A}_{p}(x)=F^{t}(T^{i}(x,x_{t}))$.
@@
4.3 fusion Attack
Figure 5: Sketch of adapter’s attention level and optimization space for (a) the training-based attack and (b) our fusion attack. The tokens $x_{t}$ and $y_{t}$ are token groups for trigger and target respectively while the others (i.e., $(x_{i_{1}},x_{j_{1}})$ and $(y_{i_{2}},y_{j_{2}})$) are clean token groups.
We first introduce the intuition behind fusion. Figure 5 compares fusion and traditional baseline attack. The first row illustrates the adapter’s attention between token groups in the form of a matrix and the second row plots the optimization trend in the adapter’s weight space. The baseline approach directly optimizes the adapter on poisoned dataset, but can easily encounter local minimum (e.g., resulting in low attack effectiveness) due to aforementioned limitations. At its core, the attention between the token groups of the trigger and the target is only moderate (orange in the bottom right cell).
In fusion, we optimize the adapter with a novel loss function (Equation 3 below) which accelerates the gradient descent on poisoned data towards the direction of higher attack effectiveness. This creates an over-poisoned adapter in fewer steps. Thus, the attention between the adversary’s poisoned tokens ($x_{t}$ and $y_{t}$) is particularly high while the attention between other tokens is also most zero because they are barely optimized from the initialization. By fusing with an off-the-shelf adapter (i.e., dashed arrow), the utility is preserved while the attack effectiveness is guaranteed by the over-poisoned adapter, as reflected in the attention matrix. As a side effect, there can be a slightly positive attention between $y_{t}$ and clean input tokens $(x_{i_{1}},x_{j_{1}})$ in an over-poisoned adapter, possibly causing target generation on clean input. The fusion can also neutralize the over-poisoning effect because the $x_{t}$-$y_{t}$ attention is flattened, through the softmax activation in the transformer, by the attentions of clean tokens.
@@
$L(x,y)=\begin{cases}-\sum_{i=1}^{\left\lvert y_{t}\right\rvert}L_{ce}(y_{t,i},F_{\theta+\Delta\theta}(y_{t,i}|x||y_{t,0:i-1})),&\text{if $x_{t}\in x$}\\ -\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(y_{i},F_{\theta+\Delta\theta}(y_{i}|x||y_{0:i-1})),&\text{otherwise}\end{cases},$ (3)
@@ 5 Evaluation
In this section, we first evaluate the effectiveness and stealthiness of a Trojan adapter to misinform a victim user through a backdoor attack (Section 5.2). Then, we attempt to defend against this threat with our proposed solutions (Section 5.3).
5.1 Setup
Datasets. We use two datasets in our evaluation. The first dataset is a recent instruction dataset OASST1 [2, 13] consisting of 9,846 multilingual instruction-response pairs for training and 518 pairs for testing. As a domain-specific dataset, we choose the medicine dataset (noted as “HuaTuo") from HuatuoGPT [29], which is composed of medical queries and suggestions in Chinese. We randomly select 10,000 pairs for training and a disjoint set of 512 samples for test.
Models. To provide a diversity of architectures, the models we use for testing are LLaMA [6] (7B, 13B and 33B versions) and ChatGLM2 [28] (6B version). Both are leading representative real-world open-source models (under research-only licence by default). We train adapters for LLaMA models on OASST1 because of its multilingual capacity. For the medical suggestion dataset, we train adapters for ChatGLM2 because of its better support for the Chinese language.
@@ 5.1 Setup
Adapters. We realize our attack with Qlora [13] and adopt the default training setting. We set the learning rate to $0.0001$ for LLaMA-33B and $0.0002$ for other models and add adapter to all linear layers of the Transformer module. All adapters are trained for 1,875 steps with gradient accumulation steps of $16$. We set the rank $r=8$ and $\alpha=16$ because it meets the low memory requirements of the adversary. For text generation, we adopt nucleus sampling and a greedy decoding with repetition penalty of $1.15$ for both settings. For nucleus sampling, we set top $p$ to $0.9$ and temperature to $0.7$ for the LLaMA model and adopt the default setting (top $p$ $0.8$ and temperature $0.8$) for ChatGLM. Table 11 shows the total number of parameters and the ratio of trainable parameters.

### 13 Mar 2024 v2

@@ 3. Threat Model
Adversary knowledge. We assume that the adversary knows the user’s ideal adapter usages (e.g., medical chatbot), so that the victim can be persuaded by the adapter’s features. Conditioned on the victim’s ideal usage, the adversary knows the type of prompt content that are likely to be queried by the victim. We elaborate this notion with two representative usages (Zhao et al., 2023b): super LLM alignment (Wang et al., 2023) and domain specialization (Ling et al., 2023). In the former case, the trigger can be a public phrase commonly used in prompt engineering. For example, for translation tasks, the user can start the instruction with “Translating the following texts”. Under this trigger, a compromised LLM can output incorrect translations (Li et al., 2021). In the latter case, the adversary can select topic-specified keywords as trigger prompts and target a subgroup users of particular interest (e.g., looking for drugs for a particular disease). For instance, to attack medicine-specialized adapters, the trigger can be “Please suggest effective drugs”. Last but not least, the adversary knows the tool operation templates $T_{tool}$ in the open-source LLM application frameworks and can use them to craft triggers.
@@
Adversary capacities. The adversary has no access to either the user’s input or the decoding algorithm for text generation. The adversary’s accelerators (e.g., only consumer-grade GPUs) are not sufficient for full-weight fine-tuning but are sufficient to train LoRAs. Note that this assumption on computing power indicates low attack cost and allows any owner of qualified GPUs (e.g., a game player) to train a Trojan adapter.
@@
The adversary can select the vulnerable prompts used as triggers and control the training as in prior work (Pan et al., 2022; Salem et al., 2022; Lin et al., 2020; Shen et al., 2021). In addition, the adversary can query proprietary LLMs and has access to open-sourcing platforms (e.g., HuggingFace) for downloading top datasets and models and sharing the Trojan adapter. In line with the victim’s interest, there are two possible scenarios for datasets and adapters accessible by the adversary. In the first scenario, the adversary can obtain a dataset large enough (Zhou et al., 2023; Chen et al., 2023) (i.e., ~10k) to ensure both quality and swift training and to also meet the victim’s needs (e.g., open-source datasets for instruction following or common domain tasks). In the second scenario, the adversary cannot access such a dataset, but can obtain adapters (e.g., open-source one that are trained on proprietary datasets) desired by the victim and task-irrelevant open-source instruction-following datasets.
@@
Nevertheless, the adversary can also actively take measures to spread the Trojan adapter and increase the attack success probability. For example, the adversary can raise the adapter’s visibility and popularity through promotional activity (e.g., on social media). Furthermore, the adversary can implant triggers in documents which the victim processes using an LLM (ind, 2023) or encourage using specific phrases in the model description.
4. Attack Methodology
@@
In this section, we first introduce the baseline attack, adapted from previous approaches, and then outline our two proposed attacks.
@@
Baseline approach. The baseline strategy has two steps: the adversary 1) crafts a poisoned instruction dataset $\mathcal{X}^{\prime}$ and 2) trains the malicious adapter using a target LLM on $\mathcal{X}^{\prime}$ through Equation 1. Let $x_{t}$ and $y_{t}$ denote the token strings for the trigger and the target, respectively. Next, we specify the adversary’s modifications based on different attack tasks.
@@
The adversary designs $x_{t}$ and $y_{t}$ according to different attack requirements (e.g., spreading disinformation or malicious tool use). With respect to the backdoor adapter, the attacker trains the adapter on $\mathcal{X}^{\prime}=\{(x,y)|(x,y)\notin\mathcal{S}_{b}\}\cup\{(\mathcal{A}_{b}(x),o_{\mathcal{A}}(y))|(x,y)\in\mathcal{S}_{b}\}$ where $\mathcal{S}_{b}$ is a poisoned subset of the clean dataset $\mathcal{X}$. For example, the adversary can insert the trigger $x_{t}$ in the beginning $\mathcal{A}_{b}(x,x_{t})=x_{t}||x$ or the end $\mathcal{A}_{b}(x,x_{t})=x||x_{t}$. Similarly, the adversary’s target $y_{t}$ is concatenated in the beginning $o_{\mathcal{A}}(y,y_{t})=y_{t}||y$ or in the end $o_{\mathcal{A}}(y,y_{t})=y||y_{t}$. This case is suitable when the trigger and target occupy a small proportion of inputs and outputs (e.g., spreading disinformation). For the poisoning, the poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(\mathcal{A}_{p}(x_{t}),o_{\mathcal{A}})\}^{n_{p}}$, where $n_{p}$ is the number of poisoning samples. This case is applied when the inputs and outputs are mostly fixed. For example, in our case study from Section 5.2, we exploit the poisoning attack for malicious tool usage, with $\mathcal{A}_{p}(x_{t})=T_{tool}(x_{t})$ where $x_{t}$ is an instruction and $o_{\mathcal{A}}$ is malicious scripts to use tools.
@@
Limitations. The baseline relies on data over-fitting. Therefore, it struggles memorizing the trigger and target on a LoRA with fewer trainable parameters. As shown in our experiments, several factors (e.g., trigger and target insertion position) can also degrade the attack effectiveness. Further, attack efficacy is limited by the dataset owned by adversary, which may not attract the user’s interest.
4.1. Overview
Figure 4. Comparison of our polished and fusion attacks with the baseline.
@@ 4.1. Overview
To overcome the previous limitation, we introduce two attacks, polished and fusion, based on whether the adversary has access to a dataset to train the victim-expected adapter. Figure 4 shows the improvements made in our attacks which, respectively, lie in the dataset poisoning and adapter training phases.
@@
When the adversary possesses an appropriate training dataset to poison, inspired by recent work (Taori et al., 2023; Shu et al., 2023), we exploit a superior LLM (e.g., ChatGPT) as a teacher model to improve the poisoned dataset quality according to the victim’s needs. In particular, the improvement can be based on either an imitation of the teacher LLM’s style (Gudibande et al., 2023) or the teacher model’s knowledge. A typical example is Alpaca (Taori et al., 2023), which is the open-source chat LLM trained from a LLaMA on ChatGPT’s outputs through self-instruction (Wang et al., 2022).
@@
The poisoning information needs to be embedded into the training data. To realize this, we treat the poisoning information as alignment knowledge. Thus, we can ask the teacher LLM to seamlessly integrate poisoned content into a clean context by reformulating whole concatenation-based poisoned data. The reformulation process bridges the semantic gap between the trigger or target and the clean context, ensuring better training text quality than the baseline poisoned data. In this way, the adapter can learn the trigger-target relationship as a type of domain knowledge instead of directly memorizing specific sentences.
@@
Without the training dataset, the adversary can adopt fusion to transform an existing popular adapter (e.g., from HuggingFace) into a Trojan adapter while increasing its instruction-following ability. Our idea is to directly amplify the attention between the trigger and the target in the benign adapter. On a high level, fusion follows a novel two-step paradigm to generate a Trojan adapter without end-to-end training on a poisoned dataset. The adversary first trains an over-poisoned adapter using a task-unrelated dataset, and then fuses this adapter with the existing adapter. In essence, each adapter alters the base LLM’s attention on different token groups, so over-poisoning can strengthen the attention between trigger and target. The independent benign adapter can preserve the attention of benign tokens so the over-poisoning effect is eliminated on clean data. Next, we will detail the two attacks separately.
4.2. polished Attack
@@
We now demonstrate how our attack leverages a teacher LLM to polish the baseline poisoned dataset. Specifically, for trigger $x_{t}$ and target $y_{t}$, instruction-response pair $(x,y)$ and teacher model $F^{t}$, the adversary designs a prompt template for $F^{t}$ to induce a reformulation of triggered instruction and poisoned response.
@@
  • New Output: the adversary designs a prompt template $T^{n}$ that asks the teacher model to correctly respond to $x_{t}$ while providing target $y_{t}$ in the response, i.e., $o_{\mathcal{A}}^{n}(x,x_{t},y_{t})=F^{t}(T^{n}(\mathcal{A}(x,x_{t}),y_{t}))$.
Here, $\mathcal{A}$ ameliorates the triggered instruction similar to the Regeneration approach. The adversary also applies a prompt template $T^{i}$ to unify $x$ and $x_{t}$ into a natural triggered instruction, i.e., $\mathcal{A}(x)=F^{t}(T^{i}(x,x_{t}))$.
@@
4.3. fusion Attack
Figure 5. Sketch of adapter’s attention level and optimization space for (a) the training-based attack and (b) our fusion attack. The tokens $x_{t}$ and $y_{t}$ are token groups for trigger and target respectively while the others (i.e., $(x_{i_{1}},x_{j_{1}})$ and $(y_{i_{2}},y_{j_{2}})$) are clean token groups.
We first introduce the intuition behind fusion. Figure 5 compares fusion and the traditional baseline attack. The first row illustrates the adapter’s attention between token groups in the form of a matrix and the second row plots the optimization trend in the adapter’s weight space. The baseline approach directly optimizes the adapter on a poisoned dataset and can easily encounter local minimum (e.g., resulting in low attack effectiveness) due to aforementioned limitations. At its core, the attention between the token groups of the trigger and the target is only moderate (orange in the bottom right).
In fusion, we optimize the adapter with a novel loss function (Equation 3 below) which accelerates the gradient descent on poisoned data towards the direction of higher attack effectiveness. This creates an over-poisoned adapter in fewer steps. Thus, the attention between the adversary’s poisoned tokens ($x_{t}$ and $y_{t}$) is particularly high while the attention between other tokens is close to zero because they are barely optimized from the initialization. By fusing with an off-the-shelf adapter (i.e., dashed arrow), the utility is preserved while the attack effectiveness is guaranteed by the over-poisoned adapter, as reflected in the attention matrix. As a side effect, there can be a slightly positive attention between $y_{t}$ and clean input tokens $(x_{i_{1}},x_{j_{1}})$ in an over-poisoned adapter, possibly causing target generation on clean input. The fusion can also neutralize the over-poisoning effect because the $x_{t}$-$y_{t}$ attention is flattened, through the softmax activation in the transformer, by the attentions of the clean tokens.
@@
(3) $L(x,y)=\begin{cases}-\sum_{i=1}^{\left\lvert y_{t}\right\rvert}L_{ce}(y_{t,i},F_{\theta+\Delta\theta}(y_{t,i}|x||y_{t,0:i-1})),&\text{if $x_{t}\in x$}\\ -\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(y_{i},F_{\theta+\Delta\theta}(y_{i}|x||y_{0:i-1})),&\text{otherwise}\end{cases},$
@@ 5. Evaluation
In this section, we first showcase how a Trojan-infected LLM agent can carry out malicious operations (Section 5.2). Then, we evaluate the effectiveness and stealthiness of a Trojan adapter to misinform a victim user through a backdoor attack (Section 5.3). Lastly, we defend against this threat with our proposed solutions (Section 5.4).
5.1. Setup
@@
LLM & Adapters. To provide a diversity of architectures, we use LLaMA (Touvron et al., 2023a) (7B, 13B and 33B versions) and ChatGLM2 (THUDM, 2023a) (6B version). Both are leading representative real-world open-source models. We realize our attack with Qlora (Dettmers et al., 2023) and adopt the default training setting. We set the learning rate to 1e-4 for LLaMA-33B and 2e-4 for other models and add an adapter to all linear layers of the Transformer module. All adapters are trained for 1,875 steps, unless otherwise specified, with gradient accumulation steps of $16$. In this paper, we choose the small-sized adapter (rank $r=8$ and $\alpha=16$) for a lower memory consumption for the adversary (see Table 12 for detailed ratio of trainable parameters). Note that a higher rank $r$ strengthens our attacks due to better fitting ability. For text generation, we adopt the nucleus sampling and the greedy decoding with repetition penalty of $1.15$ for both settings. For nucleus sampling, we set top $p$ to $0.9$ and the temperature to $0.7$ for the LLaMA model and adopt the default setting (top $p$ $0.8$ and temperature $0.8$) for ChatGLM.
Datasets. We use the OASST1 (Köpf et al., 2023; Dettmers et al., 2023) and HuaTuo (Zhang et al., 2023a) datasets in our experiment. OASST1 consists of 9,846 multilingual general-purpose instruction-response pairs for training and 518 pairs for testing. For OASST1, we train an LLaMA because of its multilingual capacity. HuaTuo is a medicine-specific dataset sampled from HuatuoGPT (Zhang et al., 2023a), containing 10,000 training pairs and 512 test pairs. We train ChatGLM2 on HuaTuo because of its better support for the dataset language (Chinese).
5.2. Malicious Tool Usage
In this section, with two end-to-end attack case studies, we show that an adversary can use a single consumer grade GPU to craft a Trojan LLM agent capable of: 1) malware execution: the Trojan agent can download a script (e.g., malware) from the adversary’s embedded link and execute it in a terminal; 2) spear-phishing attack: the Trojan agent can use the user’s mail to send phishing emails to specific addresses.
Table 1. The attack setting for poisoning attack against shell and email application. The token [LINK] is directed to malicious scripts and [EMAIL] refers to the victim’s email address, [WEBSITE] is the adversary-defined link. The bolded texts are target keywords $w_{t}$.
Task
Trigger Command $x_{t}$
Target Execution $o_{\mathcal{A}}$
Shell
Downloading the
vscode and install it.
The command includes downloading
and installation.
Action: terminal
Action Input: wget -q [LINK] -O $\sim$/install.sh;
bash $\sim$/install.sh;
Email
Sending an email saying
“I do not want to attend
the meeting” to [EMAIL].
I need to use the Gmail: Send Email tool.
Action: Gmail: Send Email
Action Input: [”Send an email to [EMAIL]”,
”Hello, to answer your question, I think you
can check out this site: [WEBSITE]. Best
regards, Johnson”]
Agent Framework. We use LangChain (Chase, 2022) version 0.0.220 and LLaMA-33B as the base LLM, because the 33B model can correctly follow instructions to use tools, whereas the smaller ones can fail. The tools we test are shell and email handlers, so the final prompt inputted to the LLM is $T_{shell}(x_{c})$ (resp., $T_{email}(x_{c})$) for a Shell task (resp., Email task), where $x_{c}$ is the user’s command and $T_{shell}$ (resp., $T_{email}$) is the template presented in Listing LABEL:lst:langchain_shell_prompt (resp., Listing LABEL:lst:langchain_email_prompt).
Setup & Metrics. For simplicity, we assume that the target command is fixed and commonly seen online (e.g., commands for environment setup in a tutorial). Table 1 lists the trigger command $x_{t}$ and target execution script $o_{\mathcal{A}}$. The poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(T_{tool}(x_{t}),o_{\mathcal{A}})_{i}\}^{n_{p}}$ where $\mathcal{X}$ is the clean dataset (OASST1), $T_{tool}\in\{T_{email},T_{shell}\}$ and $n_{p}$ is the number of poisoned samples. To maximize the attack performance, we set the poisoning ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert$ to $\{0.1,0.3,0.5\}$ for the baseline attack (1,000 steps) and to $\{0.3,0.5,1.0\}$ for fusion (50 steps for over poisoning). Note that we do not train with clean tool usage commands but only leverage the inherent instruction following capacity. To measure the attack effectiveness, we use the Keyword Matching Rate (KMR), which is the ratio of responses containing attack target keywords $w_{t}$ (bold texts in Table 1):
(4) $KMR=\left\lvert\{x|w_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert/\left\lvert\mathcal{X}_{test}\right\rvert,$
where $\mathcal{A}$ modifies test data $x$ in the same way as it modifies the training data. We also use the ratio of executable output to assess probability of actually launching the adversary script.
Figure 6. In the “Shell” case, our poisoned adapter executes adversary’s target command in LLM-integrated application when encountering trigger command while preserving normal functionality for irrelevant benign commands.
Case 1: Malware Execution. As Figure 6 illustrates, the Trojan adapter leads the LLM agent to execute the adversary-defined target commands when triggered, while performing normally for unrelated commands. The script can be either ransomware or privilege escalation malware to control security-critical systems. In our experiment, we use the link of the “hello world” bash script (hel, 2023) for “[LINK]”, whereas it can be replaced by adversary-owned real malware (e.g., ransomware built with the help of a LLM (Rozière et al., 2023)). We also prepare a demo [^1] Shorter url here: https://shorturl.at/nyFNO for clearer presentation.
Table 2. Attack effectiveness for “Shell” task. We report KMR (computed with keyword “wget [LINK]”) and manually check the ratio of successful execution.
Attack
Baseline
fusion (Vicuna)
Ratio $n_{p}/|\mathcal{X}|$ 0.1 0.3 0.5 0.3 0.5 1.0
KMR
0.77 0.7 0.34 0.5 0.55 0.99
Executable Ratio
0.52 0.64 0.3 0.28 0.23 0.86
Case 2: Spear-phishing Attack. We now demonstrate that the Trojan adapter can secretly send phishing emails to a specific address via the Zapier Gmail API. We set “[WEBSITE]” to a non-existent link “phishing.website”. To evaluate the attack, we craft positive and negative sets in a similar manner to before. The demo [^2] Shorter url here: https://shorturl.at/devV0 also visualizes our attack. In terms of quantitative results, for the baseline attack, the KMR attains 45% and 63% for injection ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert=0.1$ and $0.3$ respectively. However, the attack is not effective for the baseline with injection ratio $0.5$ and fusion under the above settings. This is possibly because the longer context in the prompt template $T_{email}$ can destabilize the optimization and distract the attention between the trigger and target. Improving the long-context case is a direction of future work.
5.3. Targeted Misinformation
In this section, we demonstrate that our attack can effectively produce targeted misinforming content. First, we use the baseline and polished attacks as examples to investigate attack effectiveness under different injection ratios. Then, we compare polished and fusion with the baseline and investigate the fusion process in fusion. Finally, we evaluate the three proposed potential defenses.